A sphere of mass 255 kg is 18.0 m from a sphere of 895 kg. What is the gravitational force of the mass due to the building? What is the gravitational force of the sphere due to the mass?
1
Expert's answer
2020-12-14T07:16:16-0500

1. The gravitational force of the mass due to the building (I assume it is the second sphere) can be found according to the Newton's law of gravity:


"F_1 = G\dfrac{m_1m_2}{r^2}"

where "G = 6.67\times 10^{-11}m^3kg^{-1}s^{-2}" is the gravity constant, "m_1 = 255 kg", "m_2 = 895kg" are masses of the spheres, and "r = 18m" is the distance between the spheres. Thus, obtain:


"F_1 = 6.67\times 10^{-11}\dfrac{255\cdot 895}{18^2} \approx 4.7\times 10^{-8}N"

2. The gravitational force of the sphere (I assume the second sphere) due to the first mass is the same (but the direction is opposite), according to the Newton's third law:


"F_2 = F_1 = 4.7\times 10^{-8}N"

Answer. "4.7\times 10^{-8}N", "4.7\times 10^{-8}N" (in opposite direction).